# Exclusive $J/\psi$, $\psi(2s)$, and $e^{+}e^{-}$ pair production in Au$+$Au ultra-peripheral collisions at RHIC

### v1

@@ Abstract
Measurements of exclusive $J/\psi$, $\psi(2s)$, and electron-positron ($e^{+}e^{-}$) pair photoproduction in Au$+$Au ultra-peripheral collisions are reported by the STAR experiment at $\sqrt{s_{{}_{\mathrm{NN}}}}=200~{}\rm{GeV}$. We report several first measurements at the Relativistic Heavy-Ion Collider, which are i) $J/\psi~{}$photoproduction with large momentum transfer up to $2.2~{}\rm{(GeV/c)^{2}}$, ii) coherent $J/\psi~{}$photoproduction associated with neutron emissions from nuclear breakup, iii) the rapidity dependence of incoherent $J/\psi~{}$photoproduction, iv) the $\psi(2s)$ photoproduction cross section at mid-rapidity, and v) $e^{+}e^{-}$ pair photoproduction up to high invariant mass of 6 $\rm{GeV/c^{2}}$. For measurement ii), the coherent $J/\psi~{}$total cross section of $\gamma^{\ast}+\rm{Au}\rightarrow J/\psi+\rm{Au}$ as a function of the center-of-mass energy $W_{\rm{\gamma*N}}$ has been obtained without photon energy ambiguities. The data are quantitatively compared with the Monte Carlo models STARlight, Sartre, BeAGLE, and theoretical calculations of gluon saturation with color glass condensate, nuclear shadowing with leading twist approximation, Quantum Electrodynamics, and the Next-to-Leading Order perturbative QCD. At the photon-nucleon center-of-mass energy of 25.0 GeV, the coherent and incoherent $J/\psi~{}$cross sections of Au nuclei are found to be $71\pm 10\%$ and $36\pm 7\%$, respectively, of that of free protons. These data provide an important experimental constraint for nuclear parton distribution functions and a unique opportunity to advance the understanding of the nuclear modification effect at the top RHIC energy.
@@ I Introduction
  • Coherent $J/\psi$ production at low $p^{2}_{\rm{T}}$ ($\lesssim 0.02~{}(\rm GeV/c)^{2}$), where both nuclei stay intact; however, the nucleus can be broken up by additional soft photons via Coulomb excitation; the primary interaction is on the nucleus level.
  • Incoherent elastic $J/\psi$ production via elastic photon-nucleon scattering at intermediate $p^{2}_{\rm{T}}$ ($\sim 0.02-0.5~{}(\rm GeV/c)^{2}$), where the target nucleus may break up into fragments; the primary interaction is on the nucleon level.
  • Incoherent dissociative $J/\psi$ production with nucleon dissociation at high $p^{2}_{\rm{T}}$ ($\gtrsim 0.5~{}(\rm GeV/c)^{2}$), where the leading nucleon (the nucleon undergoes a hard scattering) breaks up.
@@
In this paper, we report measurements of both coherent and incoherent $J/\psi~{}$photoproduction in Au$+$Au UPCs at $\sqrt{s_{{}_{\rm NN}}}$ = 200 GeV. The measurements are differential in momentum transfer $p^{2}_{\mathrm{T}}~{}$and rapidity $y$, and performed for different neutron emission classes. Furthermore, we report the first measurement of $\psi(2s)$ photoproduction at RHIC and measurements of the Quantum Electrodynamics (QED) process $\gamma\gamma\rightarrow e^{+}e^{-}$ in the invariant mass $m_{\rm{ee}}$ range of 2-6 $\rm GeV/c^{2}$.
@@ II Theoretical models
  • Sartre. A Monte Carlo model for gluon saturation physics via exclusive VM photo- and electro-production in photon-nucleus collisions. The model applies nonlinear gluon evolution to calculate the scattering amplitude of a color-dipole, provided by the photon, and a target at a given impact parameter. Based on the Good-Walker paradigm, the model can predict both coherent and incoherent VM production, where the coherent process probes the average gluon density and the incoherent process is sensitive to the density fluctuation. Also known as the hot-spot model, the mechanism in incoherent VM production implements fluctuations of the parton and nucleon positions inside of a nucleon and a nucleus, respectively, their gluon densities, and their associated saturation scale $Q_{s}$. For this analysis, the comparisons to the measured data are based on Refs. [25, 26]. This model is only valid for parton momentum fraction $x_{\rm{parton}}<0.01$, while it is compared with the STAR data at $x_{\rm{parton}}\sim 0.03$. The small mismatch to the data’s kinematics needs to be taken into account when interpreting the data.
@@ IV.1 Data selection
The UPC data were collected by the STAR experiment during the 2016 Au$+$Au run, corresponding to an integrated luminosity of 13.5 $\rm nb^{-1}$ and approximately $24\times 10^{6}$ UPC $J/\psi$-triggered events. $J/\psi~{}$candidates are selected via the electron decay channel $J/\psi~{}$$\rightarrow e^{+}e^{-}$. Based on this channel, the UPC $J/\psi~{}$trigger is defined by BEMC energy deposits greater than $\approx 0.7$ GeV in back-to-back azimuthal sextants of the BEMC. The TOF is required to have a hit multiplicity in the range of 2 to 6, and the BBCs are required to have no signal. The BBC veto and upper limit on TOF multiplicity reject most hadronic Au+Au collisions.
[…]
@@ IV.2 Simulation
  • Coherent $J/\psi~{}$$\rightarrow e^{+}e^{-}$
  • Incoherent $J/\psi~{}$$\rightarrow e^{+}e^{-}$ with elastic nucleon
  • Incoherent $J/\psi~{}$$\rightarrow e^{+}e^{-}$ with nucleon dissociation
@@
  • Coherent $\psi(2s)\rightarrow$$J/\psi~{}$$+X$, followed by $J/\psi~{}$$\rightarrow e^{+}e^{-}$ (feeddown)
@@
Two improvements were made to the generated samples. First, the STARlight $p_{\rm{T}}$ distributions of pairs from coherent $J/\psi~{}$and the two-photon processes have higher $p_{\rm{T}}$ than observed in the data. The STARlight events were reweighted by a factor of $e^{-\Delta Bp^{2}_{\rm T}}$ to describe the data, with $\Delta B=165$ (GeV/c)${}^{-2}$ for coherent $J/\psi~{}$and $\Delta B=260$ (GeV/c)${}^{-2}$ for the two-photon process. Second, the detector materials that can cause bremsstrahlung are not perfectly described by the STAR GEANT simulation, and STARlight does not include radiative processes such as $J/\psi~{}$$\rightarrow e^{+}e^{-}\gamma$. To account for this, a parallel sample of each process was made, with one electron replaced with an electron plus photon adding to the same energy. The photons were generated with a bremsstrahlung energy spectrum [52].
The generated events were passed through the simulation chain of the STAR detector. Data from an zero-bias sample (triggered on colliding bunch crossing only) of events recorded by STAR were added to the simulated events to reproduce the underlying activity in the TPC during RHIC operation. The output of this was passed though the same reconstruction algorithms as used for the data.
@@ IV.3 Signal extraction
[…]
Shown in Fig. 2 are the process templates from the simulation. Their sum is fit to the data $m_{\rm{ee}}$ and $p_{\rm{T}}$ distributions by $\chi^{2}$ minimization. It determines the fraction of extra radiative processes; the result is sufficient to account for the extra radiative effects and bremsstrahlung. The sums of all processes are also shown in Fig. 2, demonstrating a good description of the data. The rapidity distribution, not used for the fitting, demonstrates the quality of the fit.
@@
The simulated distributions are also used to determine acceptance corrections. The corrections are applied bin-by-bin to the $p_{\rm{T}}$ and $m_{\rm{ee}}$ distributions. The efficiency of the TOF $\geq 2$ hits requirement in the trigger was determined using a TOF-independent trigger, such that the TOF requirement is a complete subset of this trigger; the losses due to the $\leq 6$ TOF hits and BBC vetoes were measured using a sample of zero-bias events (triggered on colliding bunch crossing only). These two factors were applied as scale factors to the final cross sections.
@@
The fit templates are also used to separate the coherent and incoherent components of $J/\psi~{}$production, as described in Section V.2. The measured distributions $d^{2}\sigma/dp^{2}_{\rm{T}}\,dy$ are integrated over a range of $p^{2}_{\rm{T}}$, and the templates fit to the differential cross section are used to extrapolate the unmeasured region at low $p_{\rm{T}}$ and subtract the contribution from the other component. This gives the differential cross sections $d\sigma/dy$ for coherent and incoherent $J/\psi~{}$production.
@@ IV.4 Neutron tagging
The left panel of Fig. 3 shows the pulse height distribution from one of the ZDCs in the selected sample. A clear single-neutron peak at $\rm ADC=50$ is evident, well separated from the peak near zero; a two-neutron peak at $\rm ADC=120$ is also visible. The inset shows the full ADC range, exhibiting an endpoint at a pulse height corresponding to $\sim 80$ neutrons. The distribution from the other ZDC is similar. A neutron is defined as having a hit with a pulse height greater than 15 ADC counts in the ZDC, shown by the dashed line in the figure.
Neutron emissions in UPCs can provide insights to the VM production mechanism and impact parameter between the two nuclei beams. Specifically, they are categorized by their pattern of neutron emission along the beamline, which is measured by the ZDC hits on either side of the central detector, labeled as $0n$ (no hit) or $Xn$ ($\geq 1$ neutron) for each ZDC. The categories are: i) $0n0n$, neither ZDC hit; ii) $0nXn$, one ZDC hit, one not hit; iii) $XnXn$, both ZDCs hit. The sum of i) to iii) is denoted as $all~{}n$. The distribution of measured hit patterns for the full data sample is shown by the uncorrected distribution in the right of Fig. 3.
Activity from other processes in the same RHIC bunch crossing as the triggered event may include ZDC hits. These will cause migrations to ZDC categories different from those of the triggered event. The migrations will be to higher ZDC multiplicity, i.e. $0n0n$ to $0nXn$ and $XnXn$, $0nXn$ to $XnXn$. Using a sample of zero-bias events the rate of overlaps was measured to be $f_{1}=3.8\%$ for a single hit in each of the ZDCs and not the other, and $f_{2}=0.4\%$ for hits in both ZDCs; the probability of no overlaps is then $f_{0}=1-2f_{1}-f_{2}=92\%$. This determines the migration between neutron categories; the possibilities, with their probabilities in parentheses, are as follows:
@@ IV.5 Coherent $J/\psi~{}$cross sections in $\gamma$+Au collisions
Figure 4: Differential cross section $d^{2}\sigma/dp^{2}_{\mathrm{T}}dy$ of $J/\psi~{}$photoproduction as a function of $p^{2}_{\mathrm{T}}~{}$($0.005<p^{2}_{\rm{T}}<2.2~{}\rm(GeV/c)^{2}$) in Au$+$Au UPCs at $\sqrt{s_{{}_{\rm{NN}}}}=200$ GeV. The rapidity of the $J/\psi~{}$is $|y|<1.0$ and averaged $W_{\mathrm{\gamma*N}}~{}$is 25.0 GeV. MC model STARlight (upper left), nuclear shadowing of leading twist approximation calculation (upper right), Sartre MC prediction (lower left), and the color-glass-condense prediction (lower right) are compared with the data, presented as lines. Statistical uncertainty is represented by the error bars, and the systematic uncertainty is denoted as boxes. Ratio between data and models is shown in the lower sub-panel of each figure. There is a systematic uncertainty of 10% from the integrated luminosity that is not shown.
@@ IV.6 Systematic uncertainties
Several factors contribute to the acceptance and efficiency corrections for pair $m_{\rm{ee}}$ and $p_{\rm{T}}$ distributions. The trigger energy efficiency was modeled with an error function. The uncertainty of trigger efficiency on cross sections was determined by varying the BEMC tower energy threshold by 0.11 GeV as the energy resolution of the BEMC, and results in cross section uncertainties $\sim 8\%$; it is greater at low $m_{\rm{ee}}$ close to the trigger threshold. The efficiency of matching tracks to BEMC energy deposits is evaluated with TOF matched electrons in the data, resumting in an uncertainty of $\sim 5\%$ on pair reconstruction efficiency. The uncertainty on weighting of STARlight to match $p_{\rm{T}}$ distributions is only significant on the steeply falling coherent $J/\psi~{}$peak near $p_{\rm{T}}\approx 0.1$ GeV/c, where the pair uncertainty is up to 15%. The uncertainty from additional radiative events in the simulation has an uncertainty of $\sim 2\%$ on the pair acceptance.
The background subtraction with fit templates described in Section IV.3 has an statistical uncertainty from the fit. It is largest in the region of $\psi(2s)$ feeddown, $0.40.2<p_{T}<$ 0.4 GeV/c, where it reaches $\sim$10for all neutron categories except for $0n0n$ where it dominates and reaches $\sim$50%. The uncertainty from background subtraction is $<2\%$ outside this $p_{T}$ region. For the $\gamma\gamma$ $m_{ee}$ distributions the J/$\psi$ region is not reported and the background subtraction outside this region is also negligible.
@@
[…]
There is an uncertainty of 10% on the luminosity measurement, resulting in a scale uncertainty of 10% on all cross sections, which is not displayed with the data. All other uncertainties are shown with the displayed data points for all results presented hereafter. All systematic uncertainty sources are added in quadrature to obtain the total systematic uncertainty, which has an average value of $\sim$13.2%. See Table 1 for a summary of uncertainty on the cross section measurements.

### v3

# STAR Collaboration

@@ Abstract
Measurements of exclusive $J/\psi$, $\psi(2s)$, and electron-positron ($e^{+}e^{-}$) pair photoproduction in Au$+$Au ultra-peripheral collisions are reported by the STAR experiment at $\sqrt{s_{{}_{\mathrm{NN}}}}=200~{}\rm{GeV}$. We report several first measurements at the Relativistic Heavy-Ion Collider, which are i) $J/\psi~{}$photoproduction with large momentum transfer up to $2.2~{}\rm{[(GeV/c)^{2}]}$, ii) coherent $J/\psi~{}$photoproduction associated with neutron emissions from nuclear breakup, iii) the rapidity dependence of incoherent $J/\psi~{}$photoproduction, iv) the $\psi(2s)$ photoproduction cross section at mid-rapidity, and v) $e^{+}e^{-}$ pair photoproduction up to high invariant mass of 6 $\rm{GeV/c^{2}}$. For measurement ii), the coherent $J/\psi~{}$total cross section of $\gamma+\rm{Au}\rightarrow J/\psi+\rm{Au}$ as a function of the center-of-mass energy $W_{\rm{\gamma N}}$ has been obtained without photon energy ambiguities. The data are quantitatively compared with the Monte Carlo models STARlight, Sartre, BeAGLE, and theoretical calculations of gluon saturation with color glass condensate, nuclear shadowing with leading twist approximation, Quantum Electrodynamics, and the Next-to-Leading Order perturbative QCD. At the photon-nucleon center-of-mass energy of 25.0 GeV, the coherent and incoherent $J/\psi~{}$cross sections of Au nuclei are found to be $71\pm 10\%$ and $36\pm 7\%$, respectively, of that of free protons. These data provide an important experimental constraint for nuclear parton distribution functions and a unique opportunity to advance the understanding of the nuclear modification effect at the top RHIC energy.
@@ I Introduction
    Coherent $J/\psi$ production at low $p^{2}_{\rm{T}}$ (≲$\lesssim 0.02~{}(\rm GeV/c)^{2}$), where both nuclei stay intact; however, the nucleus can be broken up by additional soft photons via Coulomb excitation; the primary interaction is on the nucleus level.
    Incoherent elastic $J/\psi$ production via elastic photon-nucleon scattering at intermediate $p^{2}_{\rm{T}}$ (≈$\approx 0.02-0.5~{}(\rm[GeV/c)^{2}]$), where the target nucleus may break up into fragments; the primary interaction is on the nucleon level.
    Incoherent dissociative $J/\psi$ production with nucleon dissociation at high $p^{2}_{\rm{T}}$ (≳$\gtrsim 0.5~{}\rm[(GeV/c)^{2}]$), where the leading nucleon (the nucleon undergoes a hard scattering) breaks up.
@@
In this paper, along with results reported in a short Letter [23], we report measurements of both coherent and incoherent $J/\psi~{}$photoproduction in Au$+$Au UPCs at $\sqrt{s_{{}_{\rm NN}}}$ = 200 GeV. The measurements are differential in momentum transfer $p^{2}_{\mathrm{T}}~{}$and rapidity $y$, and performed for different neutron emission classes. Furthermore, we report the first measurement of $\psi(2s)$ photoproduction at RHIC and measurements of the Quantum Electrodynamics (QED) process $\gamma\gamma\rightarrow e^{+}e^{-}$ in the invariant mass $m_{\rm{ee}}$ range of 2-6 $\rm GeV/c^{2}$.
@@ II Theoretical models
    Sartre. A Monte Carlo model for gluon saturation physics via exclusive VM photo- and electro-production in photon-nucleus collisions. The model applies nonlinear gluon evolution to calculate the scattering amplitude of a color-dipole, provided by the photon, and a target at a given impact parameter. Based on the Good-Walker paradigm, the model can predict both coherent and incoherent VM production, where the coherent process probes the average gluon density and the incoherent process is sensitive to the density fluctuation. Also known as the hot-spot model, the mechanism in incoherent VM production implements fluctuations of the parton and nucleon positions inside of a nucleon and a nucleus, respectively, their gluon densities, and their associated saturation scale $Q_{s}$. For this analysis, the comparisons to the measured data are based on Refs. [26, 27]. This model is only valid for parton momentum fraction $x_{\rm{parton}}<0.01$, while it is compared with the STAR data at $x_{\rm{parton}}\approx 0.03$. The small mismatch to the data’s kinematics needs to be taken into account when interpreting the data.
@@ IV.1 Data selection
The UPC data were collected by the STAR experiment during the 2016 Au$+$Au run, corresponding to an integrated luminosity of 13.5 $\rm nb^{-1}$ and approximately $24\times 10^{6}$ UPC $J/\psi$-triggered events. The integrated luminosity is estimated based on the ZDC and VPD coincidence rate and the known hadronic Au$+$Au collision cross section. The final luminosity is corrected to the vertex $z$ range used in this analysis.
$J/\psi~{}$candidates are selected via the electron decay channel $J/\psi~{}$→$\rightarrow e^{+}e^{-}$. Based on this channel, the UPC $J/\psi~{}$trigger is defined by BEMC energy deposits greater than ≈$\approx 0.7$ GeV in back-to-back azimuthal sextants of the BEMC. The TOF is required to have a hit multiplicity in the range of 2 to 6, and the BBCs are required to have no signal. The BBC veto and upper limit on TOF multiplicity reject most hadronic Au+Au collisions.
Offline, pairs of tracks from a vertex within 100 cm of the center of STAR are considered. The tracks must extrapolate to energy deposits in the BEMC consistent with the trigger. The tracks must have at least 15 points in the TPC to provide good momentum resolution, and at least 11 $dE/dx$ measurements to provide good particle identification, out of a possible 45. The measure $dE/dx$ information for a track was expressed as a number of standard deviations from a particle identity hypothesis $A$, $n\sigma_{A}$. A measure of quality for the hypothesis for a pair is $\chi^{2}_{AB}=(n\sigma_{A})^{2}+(n\sigma_{B})^{2}$. Tracks consistent with electron pairs were selected by requiring $\chi^{2}_{ee}<10$ and those consistent with pion pairs were rejected by requiring $\chi^{2}_{ee}<\chi^{2}_{\pi\pi}$. Events with more than 6 significant energy deposits in the BEMC were rejected, providing further elimination of hadronic Au+Au collisions. The back-to-back requirement in the trigger is inefficient for low mass $m_{\rm{ee}}$, low $p_{\rm{T}}$ pairs, thus only pairs with $m_{\rm{ee}}>2$ GeV/c2 are included in this analysis. After applying all selection criteria, the sample includes approximately $3.9\times 10^{4}$ pairs, $7.9\times 10^{3}$ of which are in the $J/\psi~{}$mass range of $3.23.0<m_{\rm{ee}}<$ 3.2 GeV/c2.
@@ IV.2 Simulation
    Coherent $J/\psi~{}$→$\rightarrow e^{+}e^{-}$
    Incoherent $J/\psi~{}$→$\rightarrow e^{+}e^{-}$ with elastic nucleon
    Incoherent $J/\psi~{}$→$\rightarrow e^{+}e^{-}$ with nucleon dissociation
@@
    Coherent $\psi(2s)\rightarrow$$J/\psi~{}$$+X$, followed by $J/\psi~{}$→$\rightarrow e^{+}e^{-}$ (feeddown)
@@
[…]
The generated events were passed through the simulation chain of the STAR detector. Data from an zero-bias sample (triggered on colliding bunch crossing only) of events recorded by STAR were added to the simulated events to reproduce the underlying activity in the TPC during RHIC operation. The output of this was passed through the same reconstruction algorithms as used for the data.
@@ IV.3 Signal extraction
Figure 2 shows the pair mass $m_{\rm{ee}}$, transverse momentum $p_{\rm{T,ee}}$, and rapidity $y_{\rm{ee}}$ distributions. The mass distribution shown considers only pairs at low $p_{\rm{T}}<0.15$ GeV/c, where the coherent $J/\psi~{}$and $\gamma\gamma\rightarrow e^{+}e^{-}$ processes dominate. The $p_{\rm{T}}$ distribution is in the $J/\psi~{}$mass range of $3.23.0<m_{\rm{ee}}<$ 3.2 GeV/c2. The rapidity distribution includes the full selected data sample and shows the bins used for further analysis: $|y_{\rm{ee}}|<0.2$, $0.50.2<|y_{\rm{ee}}|<$ 0.5, and $10.5<|y_{\rm{ee}}|<$ 1.
Shown in Fig. 2 are the process templates from the simulation. Their sum is fit to the data $m_{\rm{ee}}$ and $p_{\rm{T}}$ distributions by $\chi^{2}$ minimization. The fit is performed simultaneously on the $m_{\rm{ee}}$ and $p_{\rm{T}}$ distributions with the same underlying parameters, which are the coherent $J/\psi~{}$yield, relative contribution between coherent and incoherent $J/\psi~{}$production, incoherent $J/\psi~{}$with and without nucleon dissociation, $\psi(2s)$ decay, and QED $\gamma\gamma$ process. It determines the fraction of extra radiative processes; the result is sufficient to account for the extra radiative effects and bremsstrahlung. The sums of all processes are also shown in Fig. 2, demonstrating a good description of the data. The rapidity distribution, not used for the fitting, demonstrates the quality of the fit.
@@
The simulated distributions are also used to determine acceptance corrections. The corrections are applied bin-by-bin to the $p_{\rm{T}}$ and $m_{\rm{ee}}$ distributions. The efficiency of the TOF ≥$\geq 2$ hits requirement in the trigger was determined using a TOF-independent trigger, such that the TOF requirement is a complete subset of this trigger; the losses due to the ≤$\leq 6$ TOF hits and BBC vetoes were measured using a sample of zero-bias events (triggered on colliding bunch crossing only). These two factors, 20% and 4%, respectively, were applied as scale factors (1.25 and 1.04) to the final cross sections.
@@
The fit templates are also used to separate the coherent and incoherent components of $J/\psi~{}$production, as described in Section V.2. The measured distributions $d^{2}\sigma/dp^{2}_{\rm{T}}\,dy$ are integrated over a range of $p^{2}_{\rm{T}}$, and the templates fit to the differential cross section are used to extrapolate the unmeasured region at low $p_{\rm{T}}$ and subtract the contribution from the other component. This gives the differential cross sections $d\sigma/dy$ for coherent and incoherent $J/\psi~{}$production. Note that this analysis does not separate incoherent contributions with and without dissociation. Only the total incoherent $J/\psi~{}$production is reported.
@@ IV.4 Neutron tagging
The left panel of Fig. 3 shows the pulse height distribution from one of the ZDCs in the selected sample. A clear single-neutron peak at $\rm ADC=50$ is evident, well separated from the peak near zero; a two-neutron peak at $\rm ADC=120$ is also visible. The inset shows the full ADC range, exhibiting an endpoint at a pulse height corresponding to ≈$\approx 80$ neutrons. The distribution from the other ZDC is similar. A neutron is defined as having a hit with a pulse height greater than 15 ADC counts in the ZDC, shown by the dashed line in the figure.
Neutron emissions in UPCs can provide insights to the VM production mechanism and impact parameter between the two nuclei beams. Specifically, they are categorized by their pattern of neutron emission along the beamline, which is measured by the ZDC hits on either side of the central detector, labeled as $0n$ (no hit) or $Xn$ (≥$\geq 1$ neutron) for each ZDC. The categories are: i) $0n0n$, neither ZDC hit; ii) $0nXn$, one ZDC hit, one not hit; iii) $XnXn$, both ZDCs hit. The sum of i) to iii) is denoted as $all~{}n$. The distribution of measured hit patterns for the full data sample is shown by the uncorrected distribution in the right of Fig. 3.
Activity from other processes in the same RHIC bunch crossing as the triggered event may include ZDC hits. These will cause migrations to ZDC categories different from those of the triggered event. The migrations will be to higher ZDC multiplicity, i.e. $0n0n$ to $0nXn$ and $XnXn$, $0nXn$ to $XnXn$. Using a sample of zero-bias events, taken during the same time as the UPC $J/\psi$ trigger, the rate of overlaps was measured to be $f_{1}=3.8\%$ for a single hit in each of the ZDCs and not the other, and $f_{2}=0.4\%$ for hits in both ZDCs; the probability of no overlaps is then $f_{0}=1-2f_{1}-f_{2}=92\%$. This determines the migration between neutron categories; the possibilities, with their probabilities in parentheses, are as follows:
@@ IV.5 Coherent $J/\psi~{}$cross sections in $\gamma$+Au collisions
Figure 4: Differential cross section $d^{2}\sigma/dp^{2}_{\mathrm{T}}dy$ of $J/\psi~{}$photoproduction as a function of $p^{2}_{\mathrm{T}}~{}$($0.005<p^{2}_{\rm{T}}<2.2~{}\rm[(GeV/c)^{2}]$) in Au$+$Au UPCs at $\sqrt{s_{{}_{\rm{NN}}}}=200$ GeV. The rapidity of the $J/\psi~{}$is $|y|<1.0$ and averaged $W_{\mathrm{\gamma N}}~{}$is 25.0 GeV. MC model STARlight (upper left), nuclear shadowing of leading twist approximation calculation (upper right), Sartre MC prediction (lower left), and the color-glass-condense prediction (lower right) are compared with the data, presented as lines. Statistical uncertainty is represented by the error bars, and the systematic uncertainty is denoted as boxes. Ratio between data and models is shown in the lower sub-panel of each figure. There is a systematic uncertainty of 10% from the integrated luminosity that is not shown.
@@ IV.6 Systematic uncertainties
Several factors contribute to the acceptance and efficiency corrections for pair $m_{\rm{ee}}$ and $p_{\rm{T}}$ distributions. The trigger energy efficiency was modeled with an error function. The uncertainty of trigger efficiency on cross sections was determined by varying the BEMC tower energy threshold by 0.11 GeV as the energy resolution of the BEMC, and results in cross section uncertainties ≈$\approx 8\%$; it is greater at low $m_{\rm{ee}}$ close to the trigger threshold. The efficiency of matching tracks to BEMC energy deposits is evaluated with TOF matched electrons in the data, resulting in an uncertainty of ≈$\approx 5\%$ on pair reconstruction efficiency. The uncertainty on weighting of STARlight to match $p_{\rm{T}}$ distributions is only significant on the steeply falling coherent $J/\psi~{}$peak near $p_{\rm{T}}\approx 0.1$ GeV/c, where the pair uncertainty is up to 15%. The uncertainty from additional radiative events in the simulation has an uncertainty of ≈$\approx 2\%$ on the pair acceptance.
The background subtraction with fit templates described in Section IV.3 has a statistical uncertainty from the fit. It is largest in the region of $\psi(2s)$ feeddown, $0.40.2<p_{T}<$ 0.4 GeV/c, where it reaches $\approx$10for all neutron categories except for $0n0n$ where it dominates and reaches $\approx$50%. The uncertainty from background subtraction is $<2\%$ outside this $p_{T}$ region. For the $\gamma\gamma$ $m_{ee}$ distributions the J/$\psi$ region is not reported and the background subtraction outside this region is also negligible.
@@
There is uncertainty on the $J/\psi~{}$photoproduction cross section in $\gamma+$Au→$\rightarrow J/\psi+$Au from the photon flux used in extracting this cross section. It is estimated by varying the Au radius ±$\pm 0.5~{}\rm{fm}$, the same method as adopted in Ref. [22]. The uncertainty on the cross section is found to be up to 3.5%.
There is an uncertainty of 10% on the luminosity measurement, resulting in a scale uncertainty of 10% on all cross sections, which is not displayed with the data. This uncertainty is estimated by a special run - the Van-der-Meer scan [55], which measures the instantaneous luminosity from the beam transverse sizes and currents; this sets the calibration factor used for the luminosity measurement. Note that there is an ongoing effort at RHIC trying to improve the uncertainty on the luminosity; however, it is not available at the time of this report and may be updated when it becomes available. All other uncertainties are shown with the displayed data points for all results presented hereafter. All systematic uncertainty sources are added in quadrature to obtain the total systematic uncertainty, which has an average value of $\approx$13.2%. See Table 1 for a summary of uncertainty on the cross section measurements.